# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v1

@@ B.2.1 Disentanglement
Table 5: Disentanglement effect. We feed different input conditions to our Media2Face* models and report the lip-sync metrics for both. Orig. denotes an image condition extracted from the video that corresponds to the audio. All other rows are results for denoising with other conditions that come from 7 differenet emotional images. Media2Face with THUNDER exhibits superior lip-sync in most cases.
B.2.2 Perceptual Study
We provide a comprehensive perceptual evaluation of THUNDER. We run a perceptual study on Amazon Mechanical Turk. The participants are shown two videos side-by-side. One corresponding to THUNDER and one to a baseline. The left-right order is randomized. The participants must finish watching both videos before being allowed to rate the videos on five-point Likert scale (strong preference for left, weak preference for left, indifference, weak preference for right and strong preference for right). We ask the participants to rate three aspects of the animation - lip-sync, dynamism and realism (as can be seen in the template in Fig. 10). The participants are shown 20 comparisons, generated from test audios, which where randomly selected from the RAVDESS and TCD-TIMIT test sets. In addition to that, we repeat the first 4 comparisons at the end of the study and discard the first 4 responses. This gives the participants a few examples such that they can adjust to the task. Additionally, we include 4 catch trials, where one animation is selected from Ground Truth and the other animation is clearly wrong. We discard the participants that are wrong on more than one catch trial. The complete results of the perceptual study are given in Fig. 8 (a model with a frozen-pretrained Wav2Vec 2.0) for THUNDER-F and Fig. 9 for THUNDER-T (a model with a finetuned Wav2Vec backbone).
Figure 8: Perceptual study results of THUNDER-F against other methods. We compare THUNDER-F against methods with both trainable audio encoder (DiffPoseTalk, EMOTE, FlameFormer-T, BaselineDiffusion-T) and frozen encoders (FlameFormer-F, BaselineDiffusion-F). Remarkably, THUNDER-F outperforms nearly all other methods on lip-sync, including those with the trainable audio-encoder (denoted -T) with the exception of the original DiffPoseTalk implementation. THUNDER-F is also better (or at least on par) with other methods in terms of dynamism of expressions and realism, with the only exception being DiffPoseTalk. We attribute the performance of DiffPoseTalk to the dataset (TFHP) which is a dataset of unscripted and more natural animations. To verify this, we also compare DiffPoseTalk against GT (last row). Surprisingly, the participants preferred DiffPoseTalk over GT. This suggests that our dataset combination of in-the-lab captured videos lacks a certain degree of natural motion compared to TFHP.
Figure 9: Perceptual study results of THUNDER-T against other methods. THUNDER-T outperforms all other methods on lip-sync, dynamism and realism with the exception of DiffPoseTalk which has very similar architecture but a larger dataset of more natural motions (TFHP).
Figure 10: The perceptual study web template.
B.2.3 Extended Ablation Study
@@
Table 6: Ablation Study. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). The mid and bottom section analyze the effect of weight $w_{mts}$, with either frozen (-F) or trainable (-T) Wav2Vec2.
@@
B.2.4 Qualitative Results
@@
Figure 11: Diversity of outputs. THUNDER is capable of generating multiple plausible animations per audio. Each row of this figure contains 6 different generations denoised from different initial noise samples (the audio condition remains the same as does the depicted frame number).
@@
Supplementary Video Our supplementary video contains exhaustive set of results which demonstrate the benefits of mesh-to-speech and analysis-by-audio-synthesis. The reader is encouraged to watch the video for our qualitative evaluation.

### v2

@@ 7.2.1 Disentanglement
Table 4: Disentanglement effect. We feed different input conditions to our versions of Media2Face* models and report the lip-sync metrics for both. Orig. denotes an image condition extracted from the video that corresponds to the audio. All other rows are results for denoising with other conditions that come from 7 different emotional images. Media2Face* with THUNDER exhibits superior lip-sync in most cases.
7.2.2 Extended Ablation and Sensitivity Analysis
@@
Table 5: Ablation study and sensitivity analysis. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). While it is already present in the main paper text, we include it here for completeness. The mid and bottom sections analyze the effect of weight $w_{mts}$, with either frozen (-F) or trainable (-T) Wav2Vec2. THUNDER-T models exhibit superior lip-sync performance but it comes at the expense of reduced diversity. Furthermore, for both THUNDER-F and THUNDER-T models, increasing $w_{mts}$, while beneficial for lip animation quality, comes at an increasing expense of of diversity. Please refer to the Sup. Video for qualitative comparison.
@@
7.2.3 Perceptual Study
Figure 8: THUNDER-F perceptual study - lip-sync. The results of THUNDER-F against other methods. We compare THUNDER-F against methods with both trainable audio encoder ( FlameFormer-T, THUNDER-T w/o mesh-to-speech) and frozen encoders (FlameFormer-F, THUNDER-F w/o mesh-to-speech). THUNDER-F outperforms other methods on lip-sync, including those with the trainable audio-encoder (denoted -T).
Figure 9: THUNDER-F perceptual study - dynamism The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER THUNDER-F w/o mesh-to-speech. However, ground truth is still rated more dynamic.
Figure 10: THUNDER-F perceptual study - realism The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER-F w/o mesh-to-speech. THUNDER-T w/o mesh-to-speech is rated about as realistic as THUNDER. Remarkably, the participants have preferred THUNDER’s generation over ground truth in terms of realism..
Figure 11: THUNDER-T complete perceptual study. The results of THUNDER-T against other methods. THUNDER-T outperforms its counterpart with mesh-to-speech and the deterministic FlameFormer on lip-sync quality. Additionally, the participatns have a slight preference for THUNDER-T over ground truth. This suggests, that while THUNDER-T exhibits a degree of dynamis, the GT potential has not yet been fully exhausted.
Figure 12: THUNDER-T perceptual study - dynamism. The participants rate THUNDER-T to be more dynamic than THUNDER-T without mesh-to-speech and also more dynamic than FlameFormer-T, THUNDER-T w/o mesh-to-speech and even Ground Truth. This suggest that THUNDER-T’s realism exhausted the realism of the dataset.
Figure 13: THUNDER-T perceptual study - realism. The participants prefer the degree of realism of THUNDER-T’s outputs over FlameFormer-T
Here we provide the rest of our comprehensive perceptual evaluation of THUNDER. We run a perceptual study on Amazon Mechanical Turk. The participants are shown two videos side-by-side. One corresponding to THUNDER and one to a baseline. The left-right order is randomized. The participants must finish watching both videos before being allowed to rate the videos on a five-point Likert scale (strong preference for left, weak preference for left, indifference, weak preference for right and strong preference for right). We ask the participants to rate three aspects of the animation - lip-sync, dynamism and realism. The exact task description can be seen in the study template in Fig. 14. The participants are shown 20 comparisons, generated from test audios, which where randomly selected from the RAVDESS and TCD-TIMIT test sets. In addition to that, we repeat the first 4 comparisons at the end of the study and discard the first 4 responses. This gives the participants a few examples such that they can adjust to the task. Additionally, we include 4 catch trials, where one animation is selected from Ground Truth and the other animation is clearly wrong. We discard the participants that are wrong on more than one catch trial. The complete results of the perceptual study for THUNDER for lip-sync, dynamism and realism can be found in Figures 10, 10 and 10 respectively. The results for THUNDER-T are given in Figures 13, 13 and 13. The results validate our qualitative and quantiative findings - THUNDER models exhibit better lip-sync, realism and dynamism when compared to baselines.
Figure 14: The perceptual study web template.
7.2.4 Generalization to other datasets
To verify whether the performance of THUNDER and mesh-to-speech translates to other types of data, we run additional experimetns on TFHP (the DiffPoseTalk dataset [sun2024diffposetalk]. TFHP was reconstructed from YouTube videos The dataset contains unscripted audios, longer sequences, and the reconstructions were obtained using MICA [MICA:ECCV2022] and SPECTRE [filntisis2022visual] and look qualitatively very different compared to our EMICA reconstructions. Can THUNDER work also in this setting?
We train both mesh-to-speech and THUNDER-T on TFHP, along with training DiffPoseTalk with the official DiffPoseTalk codebase. Please note, that in order to equalize the input conditions, we trained DiffPoseTalk without identity condioning. During evaluation, we do not use the style-condition but evaluate using the audio condition only (for both THUNDER models and DiffPoseTalk). The results are reported in Tab. 6.
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$
DiffPoseTalk (w/o style cond.) 1.24 0.341 0.508 0.25 0.0388 0.0395 0.235 0.318 0.0111
THUNDER-T w/o m2s 1.1 0.332 0.518 0.226 0.0208 0.0293 0.122 0.277 0.0068
THUNDER-T 1.09 0.351 0.545 0.214 0.0255 0.0322 0.103 0.265 0.00777
Table 6: Experiment on TFHP. We compare THUNDER-T, THUNDER w/o mesh-to-speech and DiffPoseTalk (all trained on TFHP). Similarly to our other experiments, THUNDER-T results in superior lip-sync metrics, while trading off upper-face diversity.
7.2.5 Comparison to other methods.
Comparing different speech-driven animation methods is a difficult task, especially if trained on a different dataset and especially if the data was acquired in different ways (4D scans, pseudo-GT of different reconstruction methods, etc.). Having a fair apples-to-apples comparisons stemming from measurements done against the (pseudo-) ground truth is therefore difficult. Regardless, for the sake of completeness, we compare THUNDER to the official releases of recent methods that were also trained on pseudo-GT, namely EMOTE [danecek2023emote] (a deterministic SOTA) and DiffPoseTalk [sun2024diffposetalk] (stochastic SOTA). EMOTE is a FaceFormer-inspired architecture with additional input conditions for emotions and intensity, trained with a content-emotion disentanglement mechanism. It was trained on pseudo-GT reconstructions of MEAD. The pGT was acquired with the same INFERNO tracker, making the resulting metrics comparable. The macro-architecture of DiffPoseTalk’s diffusion is very similar to THUNDER and apart from the additional style conditioning input, it differs only in small details (such as noise schedule, number of denoising steps, etc.). The most important difference is that it was trained on a different dataset, and the reconstructions were produced by a combination of SPECTRE [filntisis2022visual] and MICA [MICA:ECCV2022] and hence look qualitatively very different. As such, the DiffPoseTalk metrics likely dominated by the qualitative difference, and are only listed for completeness. Please note that for this experiment, DiffPoseTalk was run without the style condition (to match the setting of THUNDER). Finally, we include a FlameFormer trained on our data, and THUNDER trained without mesh-to-speech. The results are reported in Tab. 7.
EMOTE 1.06 0.221 0.442 0.466 0.0563 0.0401 0.248 0.24 0.0141 0.104
THUNDER 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
Table 7: Quantitative comparison of THUNDER with other speech-driven avatar methods. Asterisk* indicates we have re-implemented the baseline and trained it on our dataset. (1) On lip-sync. THUNDER outperforms all baselines on lip-sync metrics (LVE, CCC, PCC, DTW). (2) On expression diversity. DiffPoseTalk exhibits higher temporal face diversity T-DIV-U and L-DIV-U, likely due to the training data reconstructed from videos in-the-wild, which contains more dynamic motions than our in-the-lab reconstructions. EMOTE exhibits high sample upper face diversity S-DIV-U and expression diversity S-DIV-EXP which is due to the random sampling of speaking/emotion/intensity styles, which while high, may not be fitting to the particular input audio. THUNDER w/o m2s and THUNDER follow next in S-DIV-U. DiffPoseTalk (the official model) performs comparatively low on lip-sync and well on both temporal and sample diversity, but the result is likely skewed by the fact that it was trained on different dataset with different types of reconstruction.
7.2.6 Qualitative Results
@@
Figure 15: Diversity of outputs. THUNDER is capable of generating multiple plausible animations per audio. Each row of this figure contains 6 different generations of the same frame denoised from different initial noise samples (the audio condition remains the same for all).
@@
Supplementary Video. Our video contains exhaustive set of results which demonstrate the benefits of mesh-to-speech and analysis-by-audio-synthesis. The reader is encouraged to watch the video for our qualitative evaluation.
8 Architecture of Baselines
Figure 16: Architecture of re-implemented baselines. (a) Our version of Media2Face [zhao2024media2face] utilized the same architecture as THUNDER but takes an extra ImabeBind feature on the input. (c) Our re-implementation of DiffPoseTalk [sun2024diffposetalk] is also based on THUNDER and takes an extra style feature from a pretrained style encoder. (c) FlameFormer is an adapted FaceFormer [fan2022faceformer]. FlameFormer predicts FLAME expression instead of full vertex space. Furthermore, FlameFormer adapts a non-autoregressive BERT-like prediction mechanism (akin to EMOTE [danecek2023emote]).
Media2Face. Our reimplementation of Media2Face is built on our diffusion architecture. In training, we provide additional input of ImageBind[imagebind_Girdhar2023] features extracted from images of the corresponding videos of the training set. Like the audio condition, this condition is dropped 20% of the time.
DiffPoseTalk. Our reimplementation of DiffPoseTalk is also built on our diffusion architecture. It is trained in two stages. In the first stage we train the contrastive style encoder as proposed by the authors [sun2024diffposetalk]. In training of DiffPoseTalk, we provide the additional style feature on the input. The style feature is extracted from the GT animation sequence. Like the audio condition, the style condition is dropped 20% of the time.
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [danecek2023emote], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
9 Data acquisition
In this section we describe the rationale behind our choice of methodology to reconstruct the 3D faces from videos.
9.1 Choice of methodology
Acquiring high quality 4D scans of sufficient scale and richness is costly, time-consuming and requires specialized hardware. Hence, recent speech-driven animation works [sun2024diffposetalk, danecek2023emote, zhao2024media2face, yang2024probabilistic] have turned to pseudo-GT recovered from videos. This choice comes with an important decision - which face reconstruction methodolgy to use. There are several options but two of the most viable options are:
1) Optimization-based 3DDM fitting with analysis-by-image-synthesis [VetterBlanz1998, Blanz2002, Blanz2003]. While this options is viable and would produce good results, this optimization-based process is rather inefficient when deployed to many hours of video.
2) Off-the-shelf face 3DMM regressors. Due to the computational intensity of optimization-based fitting and thanks to the recent advance in deep-learning based in-the-wild face reconstruction methods [Sanyal2019_RingNet, Feng2021_DECA, EMOCA:CVPR:2021, MICA:ECCV2022, filntisis2022visual, zhang2023tokenface, Retsinas_2024_CVPR_Smirk, Tewari2017, Tewari2018, Tewari2019_FML, Deng2019]. While this is still an active area of research, and no method produces results that could be considered as accurate as 4D scans, recent years have brought enough advancement so that they can be employed for speech-driven animation dataset construction.
9.2 Discussion of FLAME regressors
Since THUNDER utilizes FLAME [FLAME:SiggraphAsia2017], our discussion will focus on FLAME regressors (trackers) only. There are several systems that have been used for talking head avatar research. No tracker Each tracker produces different kinds of errors but they were used in prior works on 3D talking head avatars. Here we briefly discuss some of the most applicable candidates:
  • DECA [Feng2021_DECA] is the first FLAME-based regressor trained with the self-supervised ”analysis-by-image-synthesis” loop. However, it is limited in terms of richness of expressions and quality of lip animations.
  • EMOCA [EMOCA:CVPR:2021] is a follow-up of DECA, capable of reconstruct rich emotions thanks to its emotion-consistency loss. However it can exaggarate expressions and the lip-sync is not good enough.
  • SPECTRE [filntisis2022visual] applies a paradigm similar to EMOCA but leverages a lip-reading consistency loss instead. It produces good lip-readable animations, but unfortunately often exaggarates, is not capable of reconstructing rich emotional expressions and the disentanglement between expression and identity is rather poor.
  • EMICA [inferno2023] is a combination of the DECA, EMOCA and SPECTRE putting together the benefits of them all to achieve high quality emotions and lip-sync. Additionally, EMICA makes use of MICA [MICA:ECCV2022] for a consistent identity prediction. The system is trained jointly and publicly available.
  • SMIRK [Retsinas_2024_CVPR_Smirk] SMIRK is a recent reconstruction method which uses a neural renderer and in-the-loop neural data augmentation to produce high quality reconstructions.
9.3 Previous works
As a result of the tremendous progress of in-the-wild face reconstruction, some of the recent SOTA speech-driven animation have already turned to pseudo-GT as the primary source of data.
  • EMOTE [danecek2023emote] employed the EMICA model to reconstruct the MEAD [kaisiyuan2020mead] dataset.
  • Yang et al. [yang2024probabilistic] utilized both DECA and SPECTRE to reconstruct LRS3 [afouras2018lrs3].
  • DiffPoseTalk employed MICA and SPECTRE separately and then combined the shape predictions of MICA and expression predictions of SPECTRE to reconstruct TFHP.
  • EMOCA exaggarates expressions.
  • DECA’s lip-sync and expressions are very limited.
  • We opt for EMICA as it provides a good compromise - rich emotions, good lip animations, consistent shape identity and temporal consistency. That said, it is not artifact-free (mouth wide open correlates with raised eyebrows, etc.) but it is sufficient to use for our research.
Following EMOTE, we opt to use EMICA in this work as well.